OpenStudy (anonymous):

solve using the system of elimination: 3x - 2y = -7 -4x + y = 11

OpenStudy (anonymous):

I got x=-3 y=-1

OpenStudy (anonymous):

Can you please show how you got this ?

OpenStudy (anonymous):

3x-2y=-7 2(-4+y=11) Then you get 3x-2y=-7 -8x+2y=22 Then you add them from each other the x-x y-y and answer-answer getting -5x=15 x=-3 then you plug in x into ONE of the problem and solve for y and you will get -1

OpenStudy (anonymous):

Thanks you so much!

OpenStudy (anonymous):

any time